OpenStudy (anonymous):

A rectangular game board is composed of identical squares arranged in a rectangular array of r rows and r+1 columns. The r rows are numbered from 1 through r, and the r+1 columns are numbered from 1 through r+1. If r>10, which of the following represents the number of squares on the board that are neither in the 4th row nor in the 7th column? A. r^2 - r B. r^2 - 1 C. r^2 D. r^2 + 1 E. r^2 + r

ganeshie8 (ganeshie8):

A

ganeshie8 (ganeshie8):

total rows : r total columns : r + 1 total squares = r(r+1) ------------(1) squares in 4th row = r+1 squares in 7th column = r squares common to above = 1 total squares in 4th row, and 7th column = 2r ------- (2) so far ok ?

OpenStudy (anonymous):

- 1 from both row (will be r-1) and column (will be r) and then multiply (will be r*r-1 or r^2-r)!